OpenStudy (anonymous):

A radar is able to detect the reflected waves from an enemy airplane,after a time interval of 0.02 milliseconds.If the velocity of the waves is 3 x 10 ms^-1, calculate the distance of the plane from the radar.

OpenStudy (anonymous):

@pooja195

OpenStudy (beginnersmind):

Assuming constant velocity what's the relation between distance traveled, velocity and time? If you get the distance traveled don't forget that the wave has to travel from the radar to the airplane _and back_.

OpenStudy (anonymous):

v = 2d/t

OpenStudy (anonymous):

3*10^8 = 2d/0.02 3*10^8 = 200d/2 3*10^8 = 100d

OpenStudy (anonymous):

3*10^5 = d?

OpenStudy (beginnersmind):

Time was given in milliseconds, not seconds. You gotta convert first.

OpenStudy (beginnersmind):

BTW, what does d mean in your expression? The distance the wave travels or the distance between the radar and the airplane?

OpenStudy (anonymous):

I'll get back to this question later,Can you help me with another question?

OpenStudy (beginnersmind):

I'll try.

OpenStudy (anonymous):

A crowbar of length 1.2 m has its fulcrum at a distance of 0.2 m from the load.Calculate the mechanical advantage of the crowbar.

OpenStudy (beginnersmind):

Sorry, can't help, I don't know the exact definition of mechanical advantage. But I guess that's what they mean. The force F on the LHS has the same torque as the force 5F on the right hand side.